OpenStudy (ledah):

Simplify: (-3y)^2(2x^2)

OpenStudy (agent47):

(-3y)^2*2x^2=9y^2*2x^2=18y^2*x^2=18(xy)^2

OpenStudy (ledah):

awesome. And a bit surprised.

OpenStudy (agent47):

I like to surprise people :D

OpenStudy (ledah):

well thanks.

OpenStudy (agent47):

np